# Modeling sea ice in the marginal ice zone as a dense granular flow with rheology inferred from a discrete element model

### v1

@@ 3 Inferring the constitutive equations of the system from the DEM
4 Analysis of the inferred continuous model
This section focuses on the one dimensional system of equations presented in section 2.2, with the functions $\mu$ and $\Phi$ taking the form (3.1). In order to understand the relative importance of the different parameters involved in this system of equations, we first rewrite it in a non-dimensional manner. Then, we study the properties of solutions to the momentum equation (2.15b) and the role of plasticity in these solutions. Finally, we consider the complete system of equations, which includes the mass conservation constraint (2.16), and we demonstrate that, under certain conditions, this system has a unique solution. For simplicity, the solutions presented throughout this section result from driving the ice floes with the ocean velocity profile (2.17), although the analysis can be extended to more general ocean velocity profiles by following the same steps.
@@ 4.1 The momentum equation
In the definition of $V$, the space $H^{1}((0,1))$ denotes the Sobolev space of weakly differentiable and periodic functions on the unit interval (Adams and Fournier,, 2003). As explained in appendix B, the functional $\mathcal{J}$ is strictly convex and coercive over $V$ and therefore admits a unique minimizer. In this sense, one can state that the momentum equation (4.2a) also has a unique solution.
@@
A similar derivation to that of (4.8) from (4.7) can be found in (Glowinski et al.,, 1981, section 1.3). In (4.8), we have introduced $\sigma_{xy}^{P}$, the purely plastic component of the shear stress. Introducing this new variable allows us to reformulate (4.2a) such that the singularity at $\mathrm{d}u/\mathrm{d}y=0$ is removed. Indeed, if $\mathrm{d}u/\mathrm{d}y\neq 0$, it is clear that (4.8) is equivalent to (4.2a). In this case, we have that $|\sigma_{xy}^{P}|=\mu_{0}p$ and we say that the material has reached its plastic yield strength $\mu_{0}p$. Conversely, when $\mathrm{d}u/\mathrm{d}y=0$, (4.8a) remains well defined, unlike (4.2a). We remark that $\mathrm{d}u/\mathrm{d}y=0$ must follow from (4.8c) whenever $|\sigma_{xy}^{P}|<\mu_{0}p$ (the material has not reached its plastic yield strength). Below, in section 4.1.2, we provide further insight into the plastic component of the shear stress by computing purely plastic solutions to the momentum equation analytically.
The first order optimality condition for the minimization of $\mathcal{J}$ are a variational inequality (rather than a variational equality) because the first term to the right hand side of (4.4) (the $L^{1}$ norm) is non differentiable when $\mathrm{d}u/\mathrm{d}y=0$. We can make $\mathcal{J}$ differentiable by regularizing it as follows:
@@
Although the system (4.8) can be solved numerically by e.g. introducing a Lagrange multiplier (Glowinski et al.,, 1981), it is easier to solve equation (4.10). This is the strategy we take for solving the momentum equation and, as we explain below in section 4.2.1, the complete system (4.2). To do so, we use the finite element method (FEM) implemented in Firedrake (Ham et al.,, 2023). In particular, we approximate the velocity profile $u$ with continuous piece-wise linear functions. In figure 5, we plot solutions to (4.10) for two different values of $p$ and a range of $\Delta>0$. Convergence of the velocity profiles as $\Delta\to 0$ is clearly visible in these figures; in fact, for $\Delta\leq 10^{-2}$, the solutions become indistinguishable. We remark that, if we remove the viscous component of the rheology in the regularized equation (4.10), we essentially arrive at Hibler’s model in one dimension, given below by (6.1a).
@@
The pressure or ice strength $p$ is a fundamental variable in the continuous model; understanding its effect on $u$ is fundamental for making sense of our sea ice model. Figure 5 suggests that, for small $p$, the velocity profile $u$ approaches the ocean’s $u_{o}$ and, for large $p$, $u$ flattens and comes close to a constant-valued velocity profile. We can deduce this behavior from the functional $\mathcal{J}$. For small values of $p$,
@@ 4.1.2 Purely plastic solutions to the momentum equation
In figure 4 we can see that $\mu(I)$ roughly becomes constant for small inertial numbers $I\ll 1$, such that $\mu(I)\approx\mu_{0}$ and the flow rheology is plastic. This regime is closely related to the quasi-static regime for granular media, with the material behaving like a purely plastic flow characterized by a critical state at which plastic deformation occurs (Wood,, 1990).
@@
4.2 Solutions to the continuous model
@@
instead of (4.2b) and (4.2c). Then, we use the FEM to solve the system of equations given by the regularized momentum equation (4.10), (4.21), and the constraint for global concentration (4.2d). We approximate $u$ with continuous piece-wise linear functions and $I_{\Delta}$ and $A$ with piece-wise constant functions. Our solver is implemented in Firedrake (Ham et al.,, 2023) in such a way that the complete nonlinear system is solved with Newton’s method. For small values of $\Delta$, Newton’s method tends to fail unless a very good initial guess for the solution $(u,I_{\Delta},A,p)$ is given. For this reason, we find that solving for a sequence of decreasing values of $\Delta$, using the solution for the previous $\Delta$ as the initial guess for the next $\Delta$, yields a robust solution method.
Figure 7: Solutions $u$ and $A$ to the continuous model (4.2) with $A_{0}=0.75$ ((a),(b)) and $A_{0}=0.9$ ((c),(d)), regularized with a parameter $\Delta$ as explained in section 4.2.1. In (e) we plot the variation of the pressure $p$ with $\Delta$. We use the same parameter values as those used for figure 5.
We test the sensitivity of solutions $(u,I_{\Delta},A,p)$ to changes in $\Delta$ by solving the system for values of $\Delta$ between $10^{-3}$ and $10$ and for $A_{0}=0.75$ and $A_{0}=0.9$. The numerical results, which are plotted in figure 7, indicate that the solutions become more sensitive to $\Delta$ as $A_{0}$ increases; this is natural, since $I$ decreases with $A_{0}$ and the plasticity term becomes more important. It is also clear from these plots that, although the velocity profiles $u$ come very close to convergence for the smallest values of $\Delta$, the local concentration profiles still experience visible changes around the symmetry points $y=0$, $0.5$, and 1. In these points the shear strain drops to 0 and, by the definition of $I$, we expect $A=1$ there. However, as we argue below in section 5, when comparing the continuous model and the DEM, we consider such drastic changes in the local concentration to be artificial. This argument motivates the use of $\Delta$ not just as a numerical parameter that helps us solve the system numerically, but as a parameter that improves the model and may have a physical significance.
@@ 4.2.2 Existence and uniqueness of solutions
We conclude the analysis of the continuous model by showing that at least one solution to (2.15) must exist and, whenever $u_{o}$ is given by (2.17), is unique. To do so, we first reformulate (2.15) solely in terms of $u$ and $p$ by eliminating $A$ and $I$ as follows. Substituting (4.2c) into (4.2d) yields
@@
In section 4.1 we show that the velocity profile $u$ has two distinct limit behaviors. We find that $u$ approaches $u_{o}$ as $p\to 0$ and that $\mathrm{d}u/\mathrm{d}y$ tends to 00 as $p\to\infty$. This means that
@@
and the function $\mathcal{F}$ appears to be strictly decreasing. This is expected, because an increase in pressure is accompanied by a flattening of the sea ice velocity profile. This means that there is a unique $p$ for which (4.26) holds; since there is only one velocity profile $u$ that solves the momentum equation (4.2a), it follows that the solution to the continuous model (4.2) must be unique.
5 Comparison of the DEM with the continuous model
The continuous model (4.2) is designed with the objective of capturing the averaged behavior of the sea ice simulations carried out with the DEM. Here, we verify that, in the one dimensional setting of the steady ocean periodic problem, the continuous model is indeed capable of replicating most of the results of the DEM. Throughout this section, we use the parameters in tables 1 and 2 when solving the continuous model. The continuous system is solved as explained in section 4.2.1 with $\Delta=10^{-3}$ and a uniform mesh of 300 cells.
Figure 9: Solutions to the non-dimensional continuous system (4.2) (black lines) compared with results from the DEM (markers), with the ocean velocity in blue. For each pair of panels in the first three rows, we fix the global sea ice concentration $A_{0}$, vary the maximum ocean velocity $u_{o,max}$ for the DEM, and plot the velocities in the panels to the left ((a), (c), (e), (g), (i), and (k)) and the local concentration to the right ((b), (d), (f), (h), (j), and (l)). In panel (m), we plot the pressure $p$ in terms of $A_{0}$. Due to the normalization in terms of $[u]=u_{o,max}$, the solutions to the continuous model are indifferent to a change in $u_{o,max}$.
@@ 5.1 Variation in global concentration $A_{0}$ and ocean speed $u_{o,max}$
We first evaluate the continuous model’s accuracy in replicating the DEM results used for fitting the functions $\mu(I)$ and $\Phi(I)$ in section 3. These results are computed for the ocean velocity profile (2.17) and an ice thickness of $H=2\,$\mathrm{m}$$. We consider six global sea ice concentrations $A_{0}$ between $0.7$ and $0.95$ and four ocean velocities $u_{o,max}$ between $0.1$ and $1\,$\mathrm{m}$/$\mathrm{s}$$. For each of these cases, we run the DEM until a steady state is reached and then we extract 10 values of the sea ice velocity and concentration along the $y$-direction, as explained in section 3, and one value for the pressure $p$, averaged over the whole square patch of ocean.
Figure 9 displays the velocity and sea ice concentration profiles $u$ and $A$ and the pressure $p$ for both the continuous model and the DEM. The velocity profile $u$ and the pressure $p$ are normalized as explained in section 4 using $[u]=u_{o,max}$. A consequence of this normalization is that the non-dimensional solutions $u$, $A$, and $p$ to the continuous model (4.2) are indifferent to the value of $u_{o,max}$. For most of the DEM results, this is also the case. For each value of $A_{0}$, almost all of the normalized velocity profiles (panels (a), (c), (e), (g), (i), and (k)) and pressure points (panel (m)) appear to collapse onto a single curve or point, which is well approximated by the continuous model.
Departures from the other normalized DEM results are most visible for slow ocean currents and low concentrations. This becomes particularly clear when looking at the pressure in panel (m) when $u_{o,max}=0.1\,$\mathrm{m}$/$\mathrm{s}$$ and $A_{0}\leq 0.8$, where the pressure values from the DEM (circles) depart substantially from the prediction of the continuous model (black solid line). These mismatches signal the continuous model’s limitations to capture the DEM results for the range of regimes considered. When fitting the dilatancy function $\Phi(I)$, the largest misfit is also found for points of slow ocean currents and low concentration ($u_{o,max}=0.1\,$\mathrm{m}$/$\mathrm{s}$$ and $A_{0}\leq 0.75$), see panel (b) in figure 4. If the DEM results are to approximate a continuous model as in (4.2), we expect the equality
@@
We remark that the term $t_{ox,j}$ results from averaging the ocean drag on each ice floe, as opposed to introducing the averaged sea ice velocity into (2.5). We also note that (5.2) ignores any choice of rheology and should hold independently of our choice of functions $\mu(I)$ and $\Phi(I)$. A surprising result we find when investigating the DEM data is that the terms to the left and right hand side of the equality in (5.2) differ in orders of magnitude whenever the sea ice concentration is low and the ocean currents are slow, see panel (a) of figure 10. Conversely, for faster ocean currents and denser concentrations, these terms become approximately equal (see panels (c)-(f) in figure 10). Inertial effects are found to be negligible in all of the cases we consider; moreover, since DEM quantities are averaged along the whole length of the domain in the $x$-direction, the term corresponding to $\partial\sigma_{xx}/\partial x$, which should be considered in a two-dimensional setting, becomes zero. Therefore, this finding raises the questions of whether a fundamentally different continuous model should be used for low sea ice concentrations and slow ocean currents or if a continuous model is valid at all.
@@
The DEM’s local concentration $A$ is accurately captured by the continuous model for low sea ice concentrations (panels (b), (d), (f), and (h) in figure 9). For $A_{0}=0.90$ and 0.95 (panels (j) and (l)), the continuous model overestimates the degree of dilatancy, although the general trends are visibly similar, with regions of higher concentration around $y=0$, $1/2$, and $1$, where the strain rates are lowest. Since an increase in $\Delta$ diminishes the local variations in $A$ (see figure 7), this suggests that $\Delta$ could be adjusted to improve the fit with the data. In this case, it would enter the model as a physical parameter whose interpretation should be explored further.
@@
Figure 11: Solutions to the non-dimensional continuous system (4.2) (black lines) compared with results from the DEM (markers). For each pair of panels in the first row, we fix the sea ice thickness $H$ and vary the maximum ocean velocity $u_{o,max}$ (the ocean velocity is plotted in blue in panels (a) and (c)). In panel (e), we plot the pressure $p$ in terms of $H$. Due to the normalization in terms of $[u]=u_{o,max}$, the solutions to the continuum system are indifferent to a change in $u_{o,max}$.

### v2

@@ 3 Inferring the constitutive equations of the system from the DEM
4 Analysis of the inferred continuum model
This section focuses on the one dimensional system of equations presented in section 2.2, with the functions $\mu$ and $\Phi$ taking the form (3.1). Firstly, in section 4.1 we non-dimensionalize the system of equations to understand the relative importance of the different terms involved.
The remaining two sections explore the existence and uniqueness of solutions to this system, a regularization technique that facilitates its numerical solution, and the behavior of solutions under different limits. In particular, in section 4.2, we focus only on the momentum equation (2.15b). In section 4.2.1, we show that the momentum equation can be rewritten as a minimization problem. This allows us to establish that solutions to this equation exist and are unique, and it allows us to make sense of the plastic component (see (3.2)) in a rigorous sense (as a variational inequality). Moreover, this equivalence also motivates a regularization of the plastic term that simplifies its numerical solution considerably, as described in section 4.2.2. We end section 4.2 with an analysis of how the sea-ice velocity behaves under different limits in parameter values: in section 4.2.3, we explore the behavior of the sea-ice velocity for small and large pressures $p$; then, in section 4.2.4, we derive an analytical solution for the sea-ice velocity in the purely plastic limit. Understanding these limit solutions for the velocity is helpful for interpreting the DEM results we present in section 5. It is also useful for the analysis we present in section 4.3, which considers the complete system of equations (2.15). We begin by presenting a numerical method for solving the complete system in section 4.3.1. Then, in section 4.3.2, we sketch out a demonstration of the existence and uniqueness of solutions to the complete system.
We remark that sections 4.2.1, 4.2.2, and 4.3 are mostly concerned with questions of a mathematical and numerical nature. Although we believe these to be important topics in establishing the suitability of our model for modeling purposes, they are not required for understanding the remainder of the paper. We also note that, for simplicity, the solutions presented throughout this section result from driving the ice floes with the ocean velocity profile (2.17), although the analysis can be extended to more general ocean velocity profiles by following the same steps.
4.1 Non-dimensionalization of the system
@@ 4.2 The momentum equation
In the definition of $V$, the space $H^{1}((0,1))$ denotes the Sobolev space of weakly differentiable and periodic functions on the unit interval (Adams & Fournier, 2003). As explained in appendix B, the functional $\mathcal{J}$ is strictly convex and coercive over $V$ and therefore admits a unique minimizer. In this sense, one can state that the momentum equation (4.2a) also has a unique solution.
@@
A similar derivation to that of (4.8) from (4.7) can be found in (Glowinski et al., 1981, section 1.3). In (4.8), we have introduced $\sigma_{xy}^{P}$, the purely plastic component of the shear stress. Introducing this new variable allows us to reformulate (4.2a) such that the singularity at $\mathrm{d}u/\mathrm{d}y=0$ is removed. Indeed, if $\mathrm{d}u/\mathrm{d}y\neq 0$, it is clear that (4.8) is equivalent to (4.2a). In this case, we have that $|\sigma_{xy}^{P}|=\mu_{0}p$ and we say that the material has reached its plastic yield strength $\mu_{0}p$. Conversely, when $\mathrm{d}u/\mathrm{d}y=0$, (4.8a) remains well defined, unlike (4.2a). We remark that $\mathrm{d}u/\mathrm{d}y=0$ must follow from (4.8c) whenever $|\sigma_{xy}^{P}|<\mu_{0}p$ (the material has not reached its plastic yield strength). Below, in section 4.2.4, we provide further insight into the plastic component of the shear stress by computing purely plastic solutions to the momentum equation analytically.
4.2.2 Regularization of the plastic term to facilitate its numerical solution
The first order optimality condition for the minimization of $\mathcal{J}$ is a variational inequality (rather than a variational equality) because the first term to the right hand side of (4.4) (the $L^{1}$ norm) is non differentiable when $\mathrm{d}u/\mathrm{d}y=0$. We can make $\mathcal{J}$ differentiable by regularizing it as follows:
@@ 4.2.2 Regularization of the plastic term to facilitate its numerical solution
Although the system (4.8) can be solved numerically by e.g. introducing a Lagrange multiplier (Glowinski et al., 1981), it is easier to solve equation (4.10). This is the strategy we take for solving the momentum equation and, as we explain below in section 4.3.1, the complete system (4.2). To do so, we use the finite element method (FEM) implemented in Firedrake (Ham et al., 2023). In particular, we approximate the velocity profile $u$ with continuous piece-wise linear functions. In figure 5, we plot solutions to (4.10) for two different values of $p$ and a range of $\Delta>0$. Convergence of the velocity profiles as $\Delta\to 0$ is clearly visible in these figures; in fact, for $\Delta\leq 10^{-2}$, the solutions become indistinguishable. We remark that, if we remove the viscous component of the rheology in the regularized equation (4.10), we essentially arrive at Hibler’s model in one dimension, given below by (6.1a).
@@
4.2.3 Velocity profiles in the limit of small and large pressures
The pressure or ice strength $p$ is a fundamental variable in the continuum model; understanding its effect on $u$ is fundamental for making sense of our sea ice model. Figure 5 suggests that, for small $p$, the velocity profile $u$ approaches the ocean’s $u_{o}$ and, for large $p$, $u$ flattens and comes close to a constant-valued velocity profile. We can deduce this behavior from the functional $\mathcal{J}$. For small values of $p$,
@@ 4.2.4 Purely plastic solutions to the momentum equation
In figure 4 we can see that $\mu(I)$ roughly becomes constant for small inertial numbers $I\ll 1$, such that $\mu(I)\approx\mu_{0}$ and the flow rheology is plastic. This regime is closely related to the quasi-static regime for granular media, with the material behaving like a purely plastic flow characterized by a critical state at which plastic deformation occurs (Wood, 1990).
@@
4.3 Solutions to the complete continuum model
@@
instead of (4.2b) and (4.2c). Then, we use the FEM to solve the system of equations given by the regularized momentum equation (4.10), (4.21), and the constraint for global concentration (4.2d). We approximate $u$ with continuous piece-wise linear functions and $I_{\Delta}$ and $A$ with piece-wise constant functions. Our solver is implemented in Firedrake (Ham et al., 2023) in such a way that the complete nonlinear system is solved with Newton’s method. For small values of $\Delta$, Newton’s method tends to fail unless a very good initial guess for the solution $(u,I_{\Delta},A,p)$ is given. For this reason, we find that solving for a sequence of decreasing values of $\Delta$, using the solution for the previous $\Delta$ as the initial guess for the next $\Delta$, yields a robust solution method.
Figure 7: Solutions $u$ and $A$ to the continuum model (4.2) with $A_{0}=0.75$ ((a),(b)) and $A_{0}=0.9$ ((c),(d)), regularized with a parameter $\Delta$ as explained in section 4.3.1. In (e) we plot the variation of the pressure $p$ with $\Delta$. We use the same parameter values as those used for figure 5.
We test the sensitivity of solutions $(u,I_{\Delta},A,p)$ to changes in $\Delta$ by solving the system for values of $\Delta$ between $10^{-3}$ and $10$ and for $A_{0}=0.75$ and $A_{0}=0.9$. The numerical results, which are plotted in figure 7, indicate that the solutions become more sensitive to $\Delta$ as $A_{0}$ increases; this is natural, since $I$ decreases with $A_{0}$ and the plasticity term becomes more important. It is also clear from these plots that, although the velocity profiles $u$ come very close to convergence for the smallest values of $\Delta$, the local concentration profiles still experience visible changes around the symmetry points $y=0$, $0.5$, and 1. In these points the shear strain drops to 0 and, by the definition of $I$, we expect $A=1$ there. However, as we argue below in section 5, when comparing the continuum model and the DEM, we consider such drastic changes in the local concentration to be artificial. This argument motivates the use of $\Delta$ not just as a numerical parameter that helps us solve the system numerically, but as a parameter that improves the model and may have a physical significance.
@@ 4.3.2 Existence and uniqueness of solutions
We conclude the analysis of the continuum model by showing that at least one solution to (2.15) must exist and, whenever $u_{o}$ is given by (2.17), is unique. To do so, we first reformulate (2.15) solely in terms of $u$ and $p$ by eliminating $A$ and $I$ as follows. Substituting (4.2c) into (4.2d) yields
@@
In section 4.2.3 we show that the velocity profile $u$ has two distinct limit behaviors. We find that $u$ approaches $u_{o}$ as $p\to 0$ and that $\mathrm{d}u/\mathrm{d}y$ tends to 00 as $p\to\infty$. This means that
@@
and the function $\mathcal{F}$ appears to be strictly decreasing. This is expected, because an increase in pressure is accompanied by a flattening of the sea ice velocity profile. This means that there is a unique $p$ for which (4.26) holds; since there is only one velocity profile $u$ that solves the momentum equation (4.2a), it follows that the solution to the continuum model (4.2) must be unique.
5 Comparison of the DEM with the continuum model
The continuum model (4.2) is designed with the objective of capturing the averaged behavior of the sea ice simulations carried out with the DEM. Here, we verify that, in the one dimensional setting of the steady ocean periodic problem, the continuum model is indeed capable of replicating most of the results of the DEM. We remind the reader that the DEM solutions can be considered steady and one-dimensional only in the sense that, for spatially and temporally-averaged data, we can expect variations in time and in the $x$-direction to be negligible. At a small scale, we certainly expect the data to present variations in time and in the $x$-direction, and the vertical velocity of the sea-ice to be non-zero. Moreover, DEM computations are initialized using a random floe initialization; however, as displayed in figure 2, the (spatially and temporally-averaged) steady states that the DEM computations reach from different floe initializations are almost indistinguishable.
Throughout this section, we use the parameters in tables 1 and 2 when solving the continuum model. The continuum model is solved with the FEM as explained in section 4.3.1, setting $\Delta=10^{-3}$ and using a uniform mesh of 300 cells.
Figure 9: Solutions to the non-dimensional continuum model (4.2) (black lines) compared with results from the DEM (markers), with the ocean velocity in blue. For each pair of panels in the first three rows, we fix the global sea ice concentration $A_{0}$, vary the maximum ocean velocity $u_{o,max}$ for the DEM, and plot the velocities in the panels to the left ((a), (c), (e), (g), (i), and (k)) and the local concentration to the right ((b), (d), (f), (h), (j), and (l)). In panel (m), we plot the pressure $p$ in terms of $A_{0}$. Due to the normalization in terms of $[u]=u_{o,max}$, the solutions to the continuum model are indifferent to a change in $u_{o,max}$.
@@ 5.1 Variation in global concentration $A_{0}$ and ocean speed $u_{o,max}$
We first evaluate the continuum model’s accuracy in replicating the DEM results used for fitting the functions $\mu(I)$ and $\Phi(I)$ in section 3. These results are computed for the ocean velocity profile (2.17) and an ice thickness of $H=2\,$\mathrm{m}$$. We consider six global sea ice concentrations $A_{0}$ between $0.7$ and $0.95$ and four ocean velocities $u_{o,max}$ between $0.1$ and $1\,$\mathrm{m}$/$\mathrm{s}$$. For each of these cases, we run the DEM until a steady state is reached and then we extract 10 values of the sea ice velocity and concentration along the $y$-direction, as explained in section 3, and one value for the pressure $p$, averaged over the whole square patch of ocean.
Figure 9 displays the velocity and sea ice concentration profiles $u$ and $A$ and the pressure $p$ for both the continuum model and the DEM. The velocity profile $u$ and the pressure $p$ are normalized as explained in section 4.1 using $[u]=u_{o,max}$. A consequence of this normalization is that the non-dimensional solutions $u$, $A$, and $p$ to the continuum model (4.2) are indifferent to the value of $u_{o,max}$. For most of the DEM results, this is also the case. For each value of $A_{0}$, almost all of the normalized velocity profiles (panels (a), (c), (e), (g), (i), and (k)) and pressure points (panel (m)) appear to collapse onto a single curve or point, which is well approximated by the continuum model.
Departures from the other normalized DEM results are most visible for slow ocean currents and low concentrations. This becomes particularly clear when looking at the pressure in panel (m) when $u_{o,max}=0.1\,$\mathrm{m}$/$\mathrm{s}$$ and $A_{0}\leq 0.8$, where the pressure values from the DEM (circles) depart substantially from the prediction of the continuum model (black solid line). These mismatches signal the continuum model’s limitations to capture the DEM results for the range of regimes considered. When fitting the dilatancy function $\Phi(I)$, the largest misfit is also found for points of slow ocean currents and low concentration ($u_{o,max}=0.1\,$\mathrm{m}$/$\mathrm{s}$$ and $A_{0}\leq 0.75$), see panel (b) in figure 4. If the DEM results are to approximate a continuum model as in (4.2), we expect the equality
@@
We remark that the term $t_{ox,j}$ results from averaging the ocean drag on each ice floe, as opposed to introducing the averaged sea ice velocity into (2.5). We also note that (5.2) ignores any choice of rheology and should hold independently of our choice of functions $\mu(I)$ and $\Phi(I)$. A surprising result we find when investigating the DEM data is that the terms to the left and right hand side of the equality in (5.2) differ in orders of magnitude whenever the sea ice concentration is low and the ocean currents are slow, see panel (a) of figure 10. Conversely, for faster ocean currents and denser concentrations, these terms become approximately equal (see panels (c)-(f) in figure 10). Inertial effects are found to be negligible in all of the cases we consider; moreover, since DEM quantities are averaged along the whole length of the domain in the $x$-direction, the term corresponding to $\partial\sigma_{xx}/\partial x$, which should be considered in a two-dimensional setting, becomes zero. Therefore, this finding raises the questions of whether a fundamentally different continuum model should be used for low sea ice concentrations and slow ocean currents or if a continuum model is valid at all.
@@
The DEM’s local concentration $A$ is accurately captured by the continuum model for low sea ice concentrations (panels (b), (d), (f), and (h) in figure 9). For $A_{0}=0.90$ and 0.95 (panels (j) and (l)), the continuum model overestimates the degree of dilatancy, although the general trends are visibly similar, with regions of higher concentration around $y=0$, $1/2$, and $1$, where the strain rates are lowest. Since an increase in $\Delta$ diminishes the local variations in $A$ (see figure 7), this suggests that $\Delta$ could be adjusted to improve the fit with the data. In this case, it would enter the model as a physical parameter whose interpretation should be explored further.
@@
Figure 11: Solutions to the non-dimensional continuum model (4.2) (black lines) compared with results from the DEM (markers). For each pair of panels in the first row, we fix the sea ice thickness $H$ and vary the maximum ocean velocity $u_{o,max}$ (the ocean velocity is plotted in blue in panels (a) and (c)). In panel (e), we plot the pressure $p$ in terms of $H$. Due to the normalization in terms of $[u]=u_{o,max}$, the solutions to the continuum model are indifferent to a change in $u_{o,max}$.